Question 3 Use the triangle below to find the measure of the angle A. A = _ (Round to one decimal place!)
Answer #1

65 LA = 2 934 BL = 6 from cos-sine formula Cosa = 6²+ (²_ q? 2 ЪС 603A - (552 +(632_cc? 2x506 608A = 25+36-16 - us 60 Cob A =,